# Vulnerabilities in AI-generated Image Detection: The Challenge of Adversarial Attacks

### v3

@@ IV-B1 Comparison with Attacks against Deepfake Detectors
TABLE III: The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets. 1 means setting (1), i.e. taking an ensemble of CNNSpot, MobileNet, and EfficientNet; 2 means setting (2), i.e. taking an ensemble of CNNSpot, MobileNet and ViT; 3 means setting (3), i.e. taking an ensemble of CNNSpot, DCTA and ViT. 4 means only using CNNSpot as surrogate model.
@@
TABLE IV: Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [3].
Attack GramNet LGrad LNP UnivFD DNF Ave
MIFGSM 9.7 6.6 38.5 24.2 46.7 25.1
PGD 67.7 50.2 47.4 11.5 48.3 45.0
S2I 49.1 49.1 28.3 8.3 43.0 35.6
FPBA 87.8 50.2 39.5 16.6 49.8 48.9
TABLE V: Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
@@
Attacks against Deepfake detectors are similar to our attacks. However, Deepfake mainly manipulating or synthesizing faces, often crafted by GANs. In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models. The content includes humans, non-human entities, abstract art, landscapes, and more. These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors. For example, techniques like searching for adversarial points on the face manifold [35] or removing facial forgery traces [36, 37] may not be applicable for AIGI detectors. To demonstrate this, we compare with FakePolisher [36] and TraceEvader [37], which are both the SOTA attacks for Deepfake detection. Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images. As shown in Table II, FPBA significantly outperforms FakePolisher and TraceEvader on both AIGI datasets. Although FakePolisher performs better than TraceEvader, but the former largely reduces the image quality, which is very visible and raises suspicion (Fig. 7). We suspect this is because FakePolisher has only learned the face manifold, limiting its ability to reconstruct fine-grained details in AIGI images.
@@ IV-B2 Comparison with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [54] and SVRE [56], i.e. utilizing an ensemble of surrogate models to generate adversarial examples. To examine the impact of different surrogate combinations(CNN-based vs. Vit-based vs. Frequency-based detectors), we conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (CNNSpot, MobileNet, EfficientNet); (2) taking CNNs and ViTs as ensemble surrogates (CNNSpot, MobileNet, ViT); (3) taking CNNs, ViTs and frequency-based detectors as ensemble surrogates (CNNSpot, DCTA, ViT). Although we can also use more than one architecture for our method, we only use CNNSpot as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Table III. First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on CNNSpot still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial example only generated by CNNSpot with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
@@ IV-B3 Evaluation on Frequency-based Detectors
TABLE VI: The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-ViT.
@@
TABLE VII: The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-ViT.
@@ IV-C1 Attack Various SOTA Detection Methods
Except for detecting in spatial and frequency domian using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts. Specifically, GramNet [53] extracts the global texture representation; UnivFD [24] captures semantic feature using pre-trained CLIP [58]; LNP [22] extracts the noise pattern of images; LGrad [23] extract gradient information through a pre-trained model and DNF [28] leverages diffusion noise features, extracting them through an inverse diffusion process [2]. These significant differences among various detection methods present challenges for the transferability of adversarial examples. To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table IV. Overall, the transfer success rate is relatively low compared to the results in Table I, suggesting that the distinct characteristics of detection methods limit adversarial transferability on each other. Nevertheless, FPBA still give the best adversarial transferability.
@@ IV-C2 Attack Cross-Generator Image Detection
The benign accuracy and attack performance on unseen source data are reported in Table V and Table VI respectively. First, the attack under the white-box setting achieves almost 100% success rate. Second, the transfer success rate is positively correlated to the accuracy on unseen source data. The Midjourney, SD v1.4&v1.5 and Wukong subsets have relatively high accuracy, their corresponding transfer attack success is also relatively high. In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them. By looking closely at the accuracy/ASR on real images and fake images(Table V,Table VII), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack. Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
IV-D Attack Performance against Compression in Real Scenarios
@@
Figure 6: Attack performace on ResNet, Efficient, MobileNet and Swin against JPEG compression on LSUN(ProGAN) dataset. Adversarial samples are crafted from ResNet.
@@
In real-world scenarios, images are often compressed during propagation, which may also distort adversarial noise patterns. Therefore, ensuring the effectiveness of adversarial attacks on compressed images is crucial for simulating real-world scenarios. To this end, we perform JPEG compression with quality factors of 30, 40, 50, 60, 70, 80, 90 and 100 to the tested images. We evaluate the degradation in attack performance using four models: ResNet-50, EfficientNet, MobileNet and Swin Transformer. All adversarial samples are crafted from ResNet-50. The results are reported in Fig. 6. For white-box attacks on ResNet, JPEG compression has a negligible effect on the attack success rate across different attack methods, maintaining top-1 attack performance. For transfer-based attacks, the attack strength on victim models decreases as the quality factor decreases. However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under the real-world scenarios.
@@
TABLE VIII: Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset. Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively. We report the attack success rate(%).
@@
IV-D1 Attack Defense Models
Furthermore, we evaluate FPBA on attacking defense models. We first employ adversarial training [38] for evaluation, which is the most popular defense baseline. However, according to our preliminary experiments(reported in the supplementary material), the training process of adversarial training fails to converge, which shows adversarial training is not a reliable way for evaluating attacks on AIGI detection. Considering that AIGI detection typically [3] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [3, 12] to utilize Gaussian blurring and JPEG compression as a defense. As shown in Table VIII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
TABLE IX: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic FFHQ datasets. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model. We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
@@ IV-D1 Attack Defense Models
FFHQ (Style GAN) FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
IV-E Visual Quality Analysis
TABLE X: The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
@@ IV-E Visual Quality Analysis
To demonstrate the superior image quality achieved by our method, we conduct both qualitative and quantitative assessments. We first visualize adversarial examples generated by PGD, FakePolisher and our proposed FPBA in Fig. 7. Adversarial examples from PGD and FakePolisher show noticeable noise patterns upon zooming in. In addition, the fine-gradined details of object generated by FakePolisher are blurred, which is very obvious in the gota’s texture and outline in its adversarial images. In contrast, FPBA generate more natural-looking adversarial examples. Further, we also report the quantitative results using common metrics for image quality assessment including MSE, PSNR(db) and SSIM. As reported in Table X, FPBA outperforms other baselines across all quality assessment by a large margin. This suggests that FPBA adding adversarial noise in the frequency-domian, rather than directly in the spatial domian, is more imperceptible to observers.
IV-F Additional Performance Analysis
IV-F1 The Phenomenon of Gradient Masking in AIGI Detectors
In Table I, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate [54]. It indicates the possibility of gradient masking [63]. To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by CNNSpot on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 8 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.

### v4

@@ IV-B1 Comparison with Attacks against Deepfake Detectors
TABLE IV: The attack success rate(%) compared with ensemble attack methods on Synthetic LSUN datasets. 1 means setting (1), i.e. taking an ensemble of CNNSpot, MobileNet, and EfficientNet; 2 means setting (2), i.e. taking an ensemble of CNNSpot, MobileNet and ViT; 3 means setting (3), i.e. taking an ensemble of CNNSpot, DCTA and ViT. 4 means only using CNNSpot as surrogate model.
@@
TABLE V: Transfer-based attack against SOTA detectors on Sythentic LSUN (ProGAN) datasets. The surrogate model is chosen as CNNSpot [3].
Attack GramNet LGrad LNP UnivFD DNF FreqNet FreqMask Ave
MIFGSM 9.7 6.6 38.5 24.2 46.7 10.5 50.3 26.6
PGD 67.7 50.2 47.4 11.5 48.3 48.9 77.2 50.2
S2I 49.1 49.1 28.3 8.3 43.0 45.3 84.2 43.9
FPBA 87.8 50.2 39.5 16.6 49.8 47.1 98.4 55.6
TABLE VI: Benign accuracy of models trained on SD V1.4 and evaluated on different generated data. ‘Real/Fake’ means the accuracy(%) on evaluating real/fake testing data.
@@
Attacks against Deepfake detectors are similar to our attacks. However, Deepfake mainly manipulate or synthesize faces, often crafted by GANs. In contrast, AIGI detectors identify any type of visual content generated by diverse techniques, such as GANs, diffusion models, and autoregressive models. The content includes humans, non-human entities, abstract art, landscapes, and more. These differences make it challenging to transfer attacks from DeepFake detectors to AIGI detectors. For example, techniques like searching for adversarial points on the face manifold [36], removing facial forgery traces [37, 38] or simply updating perturbations in the frequency domain to improve attack imperceptibility [11] may not be applicable for AIGI detectors. To demonstrate this, we compare with FakePolisher [37], TraceEvader [38] and FaceAttacker [11], which are the SOTA attacks for Deepfake detection. Because FakePolisher and TraceEvader are designed for removing forgery traces, we follow their default setting to only report the attack success rate on fake images. As shown in Table III, FPBA consistently achieves the highest attack success rates across all detectors and datasets, outperforming prior Deepfake-based attacks by margins ranging from 11.5% to 87.5%. FakePolisher performs better than TraceEvader and FaceAttack (albeit still less effective than FPBA), but the former largely reduces the image quality, which is very visible and raises suspicion, as shown in Table XIV and Fig. 8. This is likely because FakePolisher is primarily trained on the face manifold, limiting its ability to reconstruct fine-grained, identity-preserving details in AIGI images.
@@ IV-B2 Comparison with Ensemble-based Attacks
Considering that FPBA attacks an ensemble of appended models, we thus compare it with state-of-the-art ensemble-based methods ENS  [33] and SVRE [34], i.e. utilizing an ensemble of surrogate models to generate adversarial examples. To examine the impact of different surrogate combinations(CNN-based vs. Vit-based vs. Frequency-based detectors), we conduct an ablation study for ENSEMBLE and SVRE to investigate their impacts in 3 settings, including (1) taking CNNs as ensemble surrogates (CNNSpot, MobileNet, EfficientNet); (2) taking CNNs and ViTs as ensemble surrogates (CNNSpot, MobileNet, ViT); (3) taking CNNs, ViTs and frequency-based detectors as ensemble surrogates (CNNSpot, DCTA, ViT). Although we can also use more than one architecture for our method, we only use CNNSpot as the surrogate architecture to verify the universal transferability across heterogeneous models. We report the results in Table IV. First, compared with setting (1), heterogeneous model ensemble in setting (2) enhances the transferability of ENSEMBLE and SVRE across ViT-based and CNN-based models, while a more comprehensive ensemble in setting (3) decreases the transferable performance except for the frequency-based detectors. We speculate that there are significant differences in the classification boundaries between frequency-based and spatial-based detectors, thus averaging the ensemble outputs might reduce the original attack strength crafted by a single surrogate. Second, FPBA solely on CNNSpot still achieves competitive results in comparison with ensemble methods. Somewhat surprisingly, the black-box results on FPBA even outperform white-box results on ENSEMBLE and SVRE in some cases. For instance, the adversarial example only generated by CNNSpot with FPBA get success rate of 76.4% on EfficientNet, which is higher than ENSEMBLE1 and SVRE1 generated by an ensemble of CNNSpot, EfficientNet and MobileNet. This demonstrates our proposed method can approximate the true posterior distribution, in which different victim models can be sampled from the posterior.
@@ IV-B3 Evaluation on Frequency-based Detectors
TABLE VII: The attack success rate (ASR) of cross-generator image detection on different generated subsets. The surrogate model is chosen as Swin-ViT.
@@
TABLE VIII: The attack success rate (ASR) is reported on real images and fake images respectively. The surrogate model is chosen as Swin-ViT.
@@ IV-C1 Attack Various SOTA Detection Methods
Except for detecting in spatial and frequency domains using different backbone models, recent AIGI detectors utilize different latent feature representations to identify synthetic artifacts. Specifically, GramNet [54] extracts the global texture representation; UnivFD [23] captures semantic feature using pre-trained CLIP [56]; LNP [21] extracts the noise pattern of images; LGrad [22] extract gradient information through a pre-trained model; DNF [27] leverages diffusion noise features, extracting them through an inverse diffusion process [2]; FreqNet [52] enhances its generalization ability through a lightweight frequency space learning network and FreqMask [53] introduces a frequency-domain mask strategy for data augmentation during training. These significant differences among various detection methods present challenges for the transferability of adversarial examples. To illustrate this, we examine the adversarial transferability of various detection methods and present the results in Table V. Overall, the transfer success rate is relatively low compared to the results in Table II, suggesting that the distinct characteristics of detection methods limit adversarial transferability on each other. Nevertheless, FPBA still achieves the best adversarial transferability.
TABLE IX: Adversarial attack on AEROBLADE and DIRE. Average precision (AP) is used to measure the attack performance. CNNSpot is used as the surrogate model.
[…]
Clean 80.0 100.0
MIFGSM 55.9 58.9
PGD 46.9 43.2
SSI 54.6 44.5
FPBA 43.1 40.8
Very recently, reconstruction error-based methods, such as DIRE [24] and AEROBLADE [25] have emerged as competitive approaches for detecting diffusion-generated images. Therefore, we adopt them as victim models in our evaluation. As both methods rely on diffusion models, their performance degrades considerably on GAN-generated data. Consequently, we assess their adversarial robustness on the Stable Diffusion subset of the GenImage dataset in Table IX. We follow [25] to report detection performance using average precision (AP). FPBA achieves the lowest average precision (AP) on both AEROBLADE (43.1%) and DIRE (40.8%), indicating that FPBA effectively disrupts the reconstruction-error patterns these detectors rely on.
@@ IV-C2 Attack Cross-Generator Image Detection
The benign accuracy and attack performance on unseen source data are reported in Table VI and Table VII respectively. First, the attack under the white-box setting achieves almost 100% success rate. Second, the transfer success rate is positively correlated to the accuracy on unseen source data. The Midjourney, SD v1.4&v1.5 and Wukong subsets have relatively high accuracy, their corresponding transfer attack success is also relatively high. In contrast, the binary classification accuracy drops to 50% on ADM and BigGAN subsets, the corresponding adversarial transferability is also limited on them. By looking closely at the accuracy/ASR on real images and fake images(Table VI,Table VIII), we find detectors fail to distinguish fake images on ADM and BigGAN, which means without a good gradient can be followed to misclassify the real image as fake label under the attack. Therefore, we suggest robustness evaluation of cross-generator image detection should be conducted on the test subset with high accuracy as evaluating on the low-accuracy subset is futile.
Figure 6: Visualization of the spectrum saliency map (average of adversarial examples generated by FPBA using CNNSpot as the surrogate model) for different AIGI detectors.
TABLE X: Transfer attack success rate on real samples (REAL ASR) and fake samples(FAKE ASR). The surrogate model is chosen as CNNSpot.
FreqMask UnivFD GramNet LGrad FreqNet
REAL ASR 96.7 31.5 76.4 14.6 5.8
FAKE ASR 100.0 1.7 99.2 85.8 88.4
ASR 98.4 16.6 87.8 50.2 47.1
IV-D Adversarial Transferability Analysis
As shown in Table II and Table V, FPBA exhibit superior adversarial transferability compared to existing attacks. To further investigate the underlying reason, we visualize the spectrum saliency map for different AIGI detectors in Fig. 6. The map is computed as the average of adversarial examples generated by FPBA using CNNSpot as the surrogate model. Notably, CNNSpot, DenseNet, and Swin ViT share highly similar frequency-sensitive areas, leading to common vulnerabilities in these domains. In contrast, ViT exhibits fewer overlapping frequency-sensitive features with CNNSpot, resulting in a comparatively lower transfer success rate.
Next, considering that the AIGI detection task often suffers from imbalanced recognition accuracy between real and fake samples (as shown in Table VI), we further investigate the attack success rates of FPBA on real images and fake images separately, as shown in Table X. Low-level feature detectors, including FreqMask, FreqNet, GramNet, LGrad, mainly rely on frequency, texture, or spectral cues. We find that these artifacts are more susceptible to adversarial perturbations, resulting in significantly higher attack success rates on fake samples. We suspect that this vulnerability arises from low-level feature detectors overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked [60]. As a result, adversarial perturbations only need to slightly disrupt these brittle and low-dimensional features to deceive the detector, making fake images significantly more vulnerable to attack. In contrast, high-level detectors like UnivFD, which rely on global semantic features, exhibit overall lower transferability. This suggests that models leveraging generalizable semantic representations are inherently more robust. Interestingly, we observe that in UnivFD, fake images are more resistant to adversarial attacks than real ones. We will leave the phenomenon to explore in future work.
TABLE XI: AT-based defenses against PGD attack.
Dataset Detectors AT-methods Clean Robust
ProGAN CNNSpot PGD-AT 50.6 49.9
TRADES 50.1 6.5
UnivFD PGD-AT 69.6 0
TRADES 95.5 0
GenImage CNNSpot PGD-AT 50.7 49.7
TRADES 55.2 9.4
UnivFD PGD-AT 49.7 0.7
TRADES 56.8 0
IV-E Evaluation against Defense Models
Evaluation against AT-based Methods
Next, we investigate the attack performance against defense models. As adversarial training (AT) is a widely used defense strategy, we first adopt AT for improving the robustness of AIGI detectors. In Table XI, we train two representative detectors—CNNSpot and UnivFD—on the ProGAN and GenImage datasets using two mainstream AT strategies: PGD-AT [31] and TRADES [61]. However, we found that the training process of AT fails to converge. Both AT methods cause a substantial drop in clean accuracy, often to around 50%, which is equivalent to random guessing. Robust accuracy against PGD attacks also remains low, in some cases collapsing to nearly 0%. This phenomenon can be attributed to the intrinsic data imbalance in AIGI detection: models tend to overfit to low-level synthetic artifacts early in training. When AT is applied, these brittle features collapse quickly, and adversarial examples easily cross the decision boundary for both real and fake classes. As a result, the model fails to learn discriminative and robust representations, producing near-random predictions on clean data and negligible robustness gains.
Attack Performance under Compression in Real Scenarios
@@ IV-E Evaluation against Defense Models
Figure 7: Attack performance on ResNet, Efficient, MobileNet and Swin against JPEG compression on Synthetic LSUN(ProGAN) dataset. Adversarial samples are crafted from ResNet.
@@
Since AT is not a reliable defense method for AIGI detection, we need to explore alternative defense strategies. Considering that AIGI detection typically [3] uses JPEG compression and Gaussian blurring as data preprocessing during training to improve their robustness, we follow the suggestions from [3, 14] to utilize Gaussian blurring and JPEG compression as a defense. In addition, images are often compressed during propagation, which may also distort adversarial noise patterns in real-world scenarios. Therefore, ensuring the effectiveness of adversarial attacks on compressed images is also crucial for simulating real-world scenarios. To this end, we perform JPEG compression with quality factors of 30, 40, 50, 60, 70, 80, 90 and 100 to the tested images. We evaluate the degradation in attack performance using four models: ResNet-50, EfficientNet, MobileNet and Swin Transformer. All adversarial samples are crafted from ResNet-50. The results are reported in Fig. 7. For white-box attacks on ResNet, JPEG compression has a negligible effect on the attack success rate across different attack methods, maintaining top-1 attack performance. For transfer-based attacks, the attack strength on victim models decreases as the quality factor decreases. However, FPBA consistently maintains the highest attack success rate, demonstrating the effectiveness of our attack method under the real-world scenarios.
Next, we follow the suggestions from [3, 14] to simultaneously utilize Gaussian blurring and JPEG compression as a defense. As shown in Table XII, existing attack methods all suffer from performance degradation as the probability of using Gaussian blurring and JPEG compression increases during training, while FPBA still contains a high attack success rate, which validates its effectiveness.
@@
TABLE XII: Defense models(Blur+JPEG) against adversarial attack on ProGAN Dataset. Blur+JPEG(0.1) and Blur+JPEG(0.5) represented that images are blurred and JPEG-ed with 10% and 50% probability respectively. We report the attack success rate(%).
@@
TABLE XIII: The attack success rate(%) on CNN-based, ViT-based and Frequency-based models on the Synthetic FFHQ datasets. “Average” was calculated as the average transfer success rate over all victim models except for the surrogate model. We mark the white-box attack results in gray, and black-box attack results are not marked with colors.
@@
FFHQ (Style GAN)
FPBA(Ours) 51.2 55.2 66.6 100 50.3 49.9 39.4 49.8 51.8
IV-F Visual Quality Analysis
TABLE XIV: The visual quality of different attack samples in terms of the average MSE, PSNR and SSIM scores.
@@ IV-F Visual Quality Analysis
To demonstrate the superior image quality achieved by our method, we conduct both qualitative and quantitative assessments. We first visualize adversarial examples generated by PGD, FakePolisher and our proposed FPBA in Fig. 8. Adversarial examples from PGD and FakePolisher show noticeable noise patterns upon zooming in. In addition, the fine-gradined details of object generated by FakePolisher are blurred, which is very obvious in the gota’s texture and outline in its adversarial images. In contrast, FPBA generate more natural-looking adversarial examples. Further, we also report the quantitative results using common metrics for image quality assessment including MSE, PSNR(db) and SSIM. As reported in Table XIV, FPBA outperforms other baselines across all quality assessment by a large margin. This suggests that FPBA adding adversarial noise in the frequency-domian, rather than directly in the spatial domian, is more imperceptible to observers.
IV-G Additional Performance Analysis
IV-G1 The Phenomenon of Gradient Masking in AIGI Detectors
In Table II, we find that the attack results of MIFGSM are similar to IFGSM, and significantly lower than PGD, in contrast to the common belief that adversarial examples generated by momentum iterative methods have higher success rate [33]. It indicates the possibility of gradient masking [62]. To investigate whether AIGI detectors exist in gradient masking, we analyze the aggregated gradient convergence properties. We take the adversarial examples crafted by CNNSpot on ProGan as an example. For each plot, we randomly sample 500 adversarial examples generated by a specific attack to compute their expected loss gradients. Each dot shown in Fig. 9 represents a component of the expected loss gradient from each image, in which there are a total of 75k loss gradient components.